OpenStudy (anonymous):

Put the equation below in the form ax2 + bx + c = 0. Enter exponents using the caret ( ^ ). For example, you would enter 4x2 as 4x^2. -x2 + 3 = (2x - 1)2

OpenStudy (ahsome):

Ok, right now, your equation is: -x2 + 3 = (2x - 1)2 Let us insert the caret sign first. Do this were there needs to be a square root Do you know where?

OpenStudy (anonymous):

yush i think. -x^2+3=(2x-1)^2

OpenStudy (ahsome):

Yes, good job. Now its: -x^2+3=(2x-1)^2 To make it more visibaly appealing, I will make it look a bit more fancy \[-x^2+3=(2x-1)^2\] Now, we know that when we square a number, we mean multiply that number by itself 2 times. So we can rewrite that as: \[-x^2+3=(2x-1)(2x-1)\] Expand the brackets

OpenStudy (ahsome):

Try it @jackelyn_1234

OpenStudy (anonymous):

expand brackets?

OpenStudy (ahsome):

Yes. See the: \((2x-1)(2x-1)\) Expand them

OpenStudy (ahsome):

Use the foil method

OpenStudy (anonymous):

ohh! first outer inner last

OpenStudy (ahsome):

Yes, EXACTLY :) Do that for the brackets

OpenStudy (anonymous):

4x^2-4x+1

OpenStudy (ahsome):

Yes, good. Now we have: \[-x^2+3=4x^2-4x+1\] Now, we wan't to move everything so all the stuff is on one side. Can you do that? Remember, when everything is moved from one side, all we have left is 0 Example: \(x+2=2\) \(x+2-2=0\)

OpenStudy (ahsome):

want*

OpenStudy (anonymous):

-5x^4+4x+4 i believe

OpenStudy (ahsome):

I don't believe so. Here, let me help you. Equation: \[-x^2+3=4x^2-4x+1\] To move the \(3\) to the right, we need to minus \(3\) to both sides, since the \(3\) is a positive number. \[-x^2+3-3=4x^2-4x+1-3\]Adding three and taking away 3 cancel each other on the left hand side. \[-x^2\cancel{+3-3}=4x^2-4x+1-3\]\[-x^2=4x^2-4x+1-3\] Do the same thing with the \(-x^2\). This time, add \(-x^2\), since it is a negative number \[-x^2=4x^2-4x+1-3\]\[-x^2+x^2=4x^2-4x+1-3+x^2\]\[\cancel{-x^2+x^2}=4x^2-4x+1-3+x^2\]\[0=4x^2-4x+1-3+x^2\]\[4x^2-4x+1-3+x^2=0\] Now simplify the left hand side. \[4x^2-4x+1-3+x^2=0\]\[5x^2-4x+1-3=0\]\[5x^2-4x-2=0\]

OpenStudy (ahsome):

Does that make sense, @jackelyn_1234?

OpenStudy (ahsome):

Hello, @jackelyn_1234?

OpenStudy (anonymous):

i apologize. my internet wifi is acting up and my internet goes off and on

OpenStudy (anonymous):

and yes it does. i see my mistake

OpenStudy (ahsome):

Ok, good :)

OpenStudy (ahsome):

We did it. It is now in general form \((ax^2+bx+c)\) \(5x^2−4x−2=0\) In less pretty writing: 5x^2−4x−2=0

OpenStudy (ahsome):

Hope that helped you, @jackelyn_1234 :)

OpenStudy (anonymous):

thank you :D

OpenStudy (ahsome):

No problem @jackelyn_1234. But man, did it take a LOT of typing to get the equations to look pretty ;) \[\begin{bmatrix} \color{red}{WORTH}& \color{blue}{IT!} \end{bmatrix}\]